Evaluate the following convolution: (Show all your work) y(t) = {u(t+2) – u(t-2)}*{u(t+4) – u(t-4)}
Answer #1

solution :

Given y(t) = {uct +2) – 46t-2)}* {uct+u)-46-4} for Etth) 2-2 tc-6 4(E)=0 for (t+h) < 2. -6ct c-2 for Sindir I-dov = (t +4+2)